Against the wind a small plane flew 715 miles in 2 hrs and 10 minutes, the return trip took only 1 hr 50 minutes. What was the speed of the wind? What was the speed of the plane in still air?
Answer by TimothyLamb(4379)

s = d/t
x = airplane speed in still air
y = wind speed
---
against the wind:
x - y = 715/(2+10/60)
---
with the wind:
x + y = 715/(1+50/60)
---
put the system of linear equations into standard form
---
x - y = 715/(2+10/60)
x + y = 715/(1+50/60)
---
x - y = 330
x + y = 390
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = airplane speed in still air = 360 mph
y = wind speed = 30 mph
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php